OpenStudy (anonymous):

** HELP IM STUCK Johann owes $9,889 on a credit card with a 23.5% interest rate. What is the monthly payment he should make to pay off this debt in 12 months, assuming he does not charge any more purchases with the card? $824.08 $932.71 $206.30 $2,524.44

OpenStudy (anonymous):

9889-23.5%=9888.765/12= [824.06375] ?

jimthompson5910 (jim_thompson5910):

use this calculator http://www.bankrate.com/calculators/mortgages/loan-calculator.aspx

jimthompson5910 (jim_thompson5910):

alternatively, you can use this formula P = L((r/n)*(1 + r/n)^(n*t))/((1 + r/n)^(n*t) - 1) where P = monthly payment L = total amount loaned r = annual interest rate (APR) n = number of times interest is compounded per year (compounding frequency) t = time in years

OpenStudy (anonymous):

I got $932.71

jimthompson5910 (jim_thompson5910):

same here the alternative is to use the formula Plug in, L = 9889 r = 0.235 (23.5% = 23.5/100 = 0.235) n = 12 (interest is compounded monthly) t = 1 to get P = L((r/n)*(1 + r/n)^(n*t))/((1 + r/n)^(n*t) - 1) P = 9889((0.235/12)*(1 + 0.235/12)^(12*1))/((1 + 0.235/12)^(12*1) - 1) P = 932.708649237424 P = 932.71